OpenStudy (anonymous):

What is the sum of the geometric sequence 8, -16, 32 . if there are 15 terms?

OpenStudy (anonymous):

a=8 r=-16/8=-2 \[S _{n}=a \frac{ r^n-1 }{ r-1 }\] put n=15 and get the result.

OpenStudy (anonymous):

I dont understand how to do that

OpenStudy (anonymous):

\[S _{15}=8\frac{ \left( -2 \right)^{15}-1 }{ -2-1 }=8\frac{ -32768-1 }{ -3 }=8\frac{ 32769 }{ 3}=8*10923=87384\]

OpenStudy (anonymous):

Another question: What is the sum of the geometric sequence 4, 12, 36 . if there are 9 terms?

OpenStudy (anonymous):

here a=4, r=12/4=3 n=9 use the above formula